OpenStudy (aliengirl615):

What is the equation of a line that is perpendicular to y= 2x+4 and passes through the point (4, 6)?

OpenStudy (blanablas):

So you know how graphing equations and slopes and stuff like that works, right?

OpenStudy (blanablas):

Just making sure

OpenStudy (aliengirl615):

I just want to get this Apex done quickly

OpenStudy (blanablas):

Ok, y = 2x + 4 has a slope of 2

OpenStudy (blanablas):

the line you need is perpendicular, or has the inverse slope of 1/2

OpenStudy (blanablas):

wait no, -2

OpenStudy (blanablas):

brain fart there

OpenStudy (blanablas):

so far we have y = -2x

OpenStudy (aliengirl615):

is it y= -1/2x+4 ?

OpenStudy (blanablas):

if it passes through 4,6, then that is a solution to the equation, i.e 6 = -2(4) + x

OpenStudy (blanablas):

6 = -8 + x

OpenStudy (blanablas):

14 = x

OpenStudy (aliengirl615):

i got it! thanks

OpenStudy (blanablas):

so the equation is y=-2x +14 i'm sure

OpenStudy (blanablas):

wait checking my work

OpenStudy (aliengirl615):

no it was Y= 1/2x+8 i think

OpenStudy (blanablas):

is this an essay or multi-choice question?

OpenStudy (blanablas):

also you are right it is -1/2

OpenStudy (blanablas):

and if it passes through 4,6 then yeah it is y=-1/2x + 8